OpenStudy (anonymous):

How do I convert x(t)=cos(t)+sin(t-pi/6) to the form Acos(wt-phi)?

OpenStudy (anonymous):

This is more a math question

OpenStudy (anonymous):

yea but.. its related to waves.. its an equation for travelling wave!. so you have \[x(t) = \cos (t) + \sin(t- \pi/6)\] Since you want the final answer in Cos form.. convert the sin function to cos function such that you would get a positive conversion.. (using allied angles ) next apply Cos(c) + cos(d) formula and gwalla.. so can you do those steps.. if you get stuck ll help you out!